Are the following lines parallel, perpendicular, or neither?
L1 with equation x – 2y = 10
L2 with equation 2x + y = 2

Answer by checkley71(8403)

LINES ARE PERPENDICULAR X-2Y=10 OR 2Y=X-10 OR Y=1/2X-5
2X+Y=2 OR Y=-2X+2 THUS THE TWO SLOPES ARE NEGATIVE RECIPRICALS OF EACH OTHER THUS THEY ARE PERPENDICULAR.